OpenStudy (anonymous):

The girls in Karen’s family own a total of 46 T-shirts. The oldest, Heidi, has 18 T-shirts. Karen has six times as many T-shirts as her little sister, June. How many T-shirts does Karen own? The girls in Karen’s family own a total of 46 T-shirts. The oldest, Heidi, has 18 T-shirts. Karen has six times as many T-shirts as her little sister, June. How many T-shirts does Karen own? @Mathematics

OpenStudy (anonymous):

24

OpenStudy (anonymous):

how do you do the problem?

OpenStudy (anonymous):

if they are three sisters

OpenStudy (anonymous):

Karen and Heidi are the only sisters?

OpenStudy (anonymous):

18 + 6x + x = 46 7x = 46 - 18 x = 28/7 x = 4 therefore 6*4 = 24

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

24